the sum of two numbers is 141. One is 37 more than the other. find the numbers.
Answer by praseenakos@yahoo.com(507)

QUESTION:

the sum of two numbers is 141. One is 37 more than the other. find the numbers.

ANSWER:


It is given that....the sum of two numbers is 141.



Also it is given that...One is 37 more than the other. find the numbers.


Assume that one number is x.

then the other number is x + 37


Their sum is given as 141.



So we can write,

x + x + 37 = 141


==> 2x + 37 = 141



Subtract 37 from both sides.......



==> 2x + 37 - 37 = 141 - 37



==> 2x = 104



Divide both sides of the equation by 2


==> 2x/2 = 104/2



==> x = 52


So one number is 52



Other number is x + 37 = 52 + 37 = 89




while checking the answer, you can see that sum of 89 and 52 is 141.


Hope you found the explanation useful.




Regards.


Praseena.